OpenStudy (anonymous):

prove the identity \[\left( \cos x \over 1 +\sin x \right)=\left( 1-\sin x \over \cos x \right)\]

OpenStudy (anonymous):

LHS/RHS=1

OpenStudy (anonymous):

because cos x cos x = 1- sin x sin x

OpenStudy (mr.math):

By multiplying the left hand side by 1-sinx, we get \[{\cos{x}(1-\sin{x}) \over 1-\sin^2{x}}={\cos(x)(1-\sin(x) \over \cos^2(x)}={1-\sin{x} \over \cos{x}}\]

OpenStudy (mr.math):

Which is what we have on the right hand side obviously.

OpenStudy (anonymous):

that was simple. thanks.

OpenStudy (mr.math):

It is simple! :)

OpenStudy (anonymous):

Yifan you help me so much! and I really appreciate it, but many of times I don't understand the way you teach.. I am sorry. :S

OpenStudy (mr.math):

His argument is not that difficult, he's using that if a=b\(\ne\)0, then a/b=1.

OpenStudy (anonymous):

actually i should use LHS-RHS=0 to prove that